One square foot of farm land requires four seeds. At this rate, how many seeds will be required to plant ten square yards?
Answer by jim_thompson5910(28504)

1 square yard = 9 square feet

So 1/9 of a square yard requires four seeds, which means we have the idea that

1/9 sq yd = 4 seeds

Now multiply both sides by 90 to get

90*(1/9) sq yd = 90*4 seeds

90/9 sq yd = 360 seeds

10 sq yd = 360 seeds


So 360 seeds are needed to plant 10 square yards.